OpenStudy (anonymous):

find the third term of an arithmetic sequence with t2=9/2 and t5=6

OpenStudy (amistre64):

there is a generic formula for these things: an = a1 + d(n-1)

OpenStudy (amistre64):

as such we can determine the missing parts with what weve got assume t2 = t1 and t5 = t4 ; im used to using as and not ts, but that doesnt matter

OpenStudy (amistre64):

as such: t4 = t1 + d(3-1) t4 = 9/2 + d(3-1) 6 = 9/2 + 2d ; solve for d

OpenStudy (amistre64):

not 3-1; 4-1 :) t4 = 9/2 + d(4-1) 6 = 9/2 + 3d ; solve for d

OpenStudy (anonymous):

i got 4,5 but all my possible answers are singe digits

OpenStudy (anonymous):

single*

OpenStudy (anonymous):

4.5*

OpenStudy (amistre64):

well 6 - 9/2 = 3d 3/2 = 3d 1/2 = d t2 + 1/2 = t3 9/2 + 1/2 = 10/2 = 5

OpenStudy (anonymous):

hm.... i calculated wrong...

OpenStudy (anonymous):

what is the 5th term of arithmetic sequence if t2=-5 and t6=7

OpenStudy (amistre64):

its the same concept, that is the beauty of generality; it works for any specific that you need to apply it to

OpenStudy (amistre64):

hmm, it looks like we could also that arithmetic mean to do this with as well t3 t4 t5 give us 3 terms to find; so difference divided by 4 7--5 = 12/4 = 3 for a common difference t6 - d = t5 7-3 = 4 = t5, if i did that right

OpenStudy (anonymous):

i think so. the answer is 4 then? cause that would fit my answer options

OpenStudy (amistre64):

give it a shot:)

OpenStudy (anonymous):

the only ones i got wrong were the ones i either did ask here or guessed to finish... lol, but i passes that one!

OpenStudy (anonymous):

what are the first four terms of the geometric sequence with a t1=2 and tn=3tn-1